# Learning to Parallel: Accelerating Diffusion Large Language Models via Adaptive Parallel Decoding

### v1

@@ 1 Introduction
    We propose a novel and adaptive framework, Learn2PD that predicts which tokens have been correctly decoded, approximating the oracle Extremely Greedy Parallel Decoding strategy.
@@ 2.2 Accelerate Diffusion-based large language models
Followed by mature diffusion large language models, their acceleration methods are also under development. Concretely, dllm-Cache [Liu et al., 2025] proposes a training-free, adaptive caching framework that performs long-interval prompt caching and short-interval, value-similarity–guided partial response updates. Fast-dLLM [Wu et al., 2025] introduces block-wise approximate KV caching and a confidence-aware parallel decoding rule that only decodes tokens whose marginal confidence exceeds a threshold. Hu et al. [2025] propose FreeCache to approximate KV states by reusing stable prompt/block activations across steps. They also introduce Guided Diffusion to decide which tokens to unmask each step without retraining. SlowFast-Sampling [Wei et al., 2025] proposes a dynamic two-stage sampler that alternates a cautious exploratory phase with a fast phase that aggressively decodes high-confidence tokens within that span. Prophet [Li et al., 2025] monitors the top-2 logit gap and commits all remaining tokens in one shot via early-commit decoding once it is sufficiently confident. However, these acceleration methods are often static and lack flexibility. To address this, we propose Learn2PD, a novel dynamic and parallel decoding method that achieves more efficient inference acceleration by reducing the unnecessary and repetitive decoding steps. Moreover, we also introduce EoTP to avoid redundant decoding when the answer does not span the full generation length.
@@ 3.3 End-of-Text Prediction
Figure 5: Schematic of the End-of-Text Prediction Policy. During the inference process, upon detection of an [EoT] token in a decoded block, all subsequent tokens are assigned with [EoT], and the inference is halted immediately.
Besides the methods mentioned earlier, we observed that when the generation length of a diffusion large language model is increased to 1024, the generation time rises significantly for the same question compared to a length of 256, even though the final answer length remains unchanged. According to the analysis of the generated output, we find that the extra length is filled with the [EoT] token, and the additional decoding time is spent repeatedly decoding the [EoT] token. Based on this, we propose the End-of-Text Prediction (EoTP) approach: whenever the last decoded token is [EoT], the model would terminate the decoding process immediately and return the response. Therefore, we update the inference process to handle the long generation length challenge in Appendix A. We illustrate the details in Figure 5. Our analysis shows that 89.59% of computational cost comes from decoding padding tokens after [EoT]. EoTP eliminates this overhead by detecting sequence completion when all non-[MASK] positions have unmasked. The experiments and relevant analysis are in Appendix B.

### v2

# Learning to Parallel: Accelerating Diffusion Large Language Models via Learnable Parallel Decoding

@@ 1 Introduction
    We propose a learnable and adaptive framework, Learn2PD that predicts which tokens have been correctly decoded, approximating the oracle Extremely Greedy Parallel Decoding strategy.
@@ 2.2 Accelerate Diffusion-based large language models
Followed by mature diffusion large language models, their acceleration methods are also under development. Concretely, dllm-Cache [Liu et al., 2025] proposes a training-free, adaptive caching framework that performs long-interval prompt caching and short-interval, value-similarity–guided partial response updates. Fast-dLLM [Wu et al., 2025] introduces block-wise approximate KV caching and a confidence-aware parallel decoding rule that only decodes tokens whose marginal confidence exceeds a threshold. Hu et al. [2025] propose FreeCache to approximate KV states by reusing stable prompt/block activations across steps. They also introduce Guided Diffusion to decide which tokens to unmask each step without retraining. SlowFast-Sampling [Wei et al., 2025] proposes a dynamic two-stage sampler that alternates a cautious exploratory phase with a fast phase that aggressively decodes high-confidence tokens within that span. Prophet [Li et al., 2025] monitors the top-2 logit gap and commits all remaining tokens in one shot via early-commit decoding once it is sufficiently confident. These acceleration methods are often static and lack flexibility. APD [Israel et al., 2025] employs a small autoregressive model to dynamically control the number of tokens generated in parallel. But they fix the dLLM decoding process to proceed left-to-right, which prevents them from fully exploiting the parallelization advantages of dLLMs. To address this, we propose Learn2PD, a novel dynamic and parallel decoding method that achieves more efficient inference acceleration by reducing the unnecessary and repetitive decoding steps. Moreover, we also introduce EoTP to avoid redundant decoding when the answer does not span the full generation length.
@@ 3.3 End-of-Text Prediction
Figure 5: Schematic of the End-of-Text Prediction Policy. During the inference process, upon detection of an [EoT] token in a decoded block, all subsequent tokens are discarded.
Besides the methods mentioned earlier, we observed that when the generation length of a diffusion large language model is increased to 1024, the generation time rises significantly for the same question compared to a length of 256, even though the final answer length remains unchanged. According to the analysis of the generated output, we find that the extra length is filled with the [EoT] token, and the additional decoding time is spent repeatedly decoding the [EoT] token. Based on this, we propose the End-of-Text Prediction (EoTP) approach: at each step, when a [EoT] token is unmasked, discard tokens in subsequent positions. The shortened sequence is then used as the input of next step. Therefore, we update the inference process to handle the long-generation-length challenge in Appendix A. We illustrate the details in Figure 5. Our analysis shows that 89.59% of computational cost comes from decoding padding tokens after [EoT]. EoTP yields substantial computational savings by dynamically reducing the effective input length throughout the diffusion process. The experiments and relevant analysis are in Appendix B.